OpenStudy (anonymous):

James added 3 and two-thirds plus 4 and one-eighth. His work is shown here. What did James do right? What did he do wrong? What should he have done? three and two-thirds plus 4 and one-eighth equals 7 and three-elevenths

OpenStudy (anonymous):

@Chineseboy15

OpenStudy (anonymous):

@angel12310 please come help me

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[3\frac{ 2 }{ 3 }+4\frac{ 1 }{ 8 }\] First make the fractions have a common denominator. \[3\frac{ 2 }{ 3 }+4\frac{ 1 }{ 8 }\] \[ =\frac{ 11 }{ 3 }+\frac{ 33 }{ 8 }\] \[=\frac{ 11\times8 }{ 3\times8 }+\frac{ 33\times3 }{ 8\times3 }\] \[=\frac{ 88 }{ 24 }+\frac{ 99 }{ 24 }\] \[=\frac{ 187 }{ 24 }\] \[=7\frac{ 19 }{ 24 }\] James do wrong.

OpenStudy (anonymous):

I hope my reply does not come so late. It take me a long time.

OpenStudy (anonymous):

3 2/3 + 4 1/8 11/3 + 33/8 = ((11 × 8) + (33 × 3)) / (3 × 8) = (88 + 99) / 24 = 187/24 = 187/24 = 7 19/24

OpenStudy (anonymous):

it is ok, I understand :)

OpenStudy (anonymous):

wow! hold on a sec I gotta read this!

OpenStudy (anonymous):

whatever Bye

OpenStudy (anonymous):

Can you understand?

OpenStudy (anonymous):

no, I dont

OpenStudy (anonymous):

can u explain it to me?

OpenStudy (anonymous):

\[3\frac{ 2 }{ 3 }=\frac{ 11 }{ 3 }, 4\frac{ 1 }{ 8 }=\frac{ 33 }{ 8 }\], is it?

OpenStudy (anonymous):

hmm...

OpenStudy (anonymous):

3 2/3 + 4 1/8 then change them into an improper fraction: 11/3 + 33/8 = ((11 × 8) + (33 × 3)) / (3 × 8) = (88 + 99) / 24 = 187/24 = 187/24 = 7 19/24

OpenStudy (anonymous):

angel, can u in words?

OpenStudy (anonymous):

ok so the answer is?

OpenStudy (anonymous):

Well, if we want to add them together, we need to make two fractions have same denominators.

OpenStudy (anonymous):

I has to write it you know

OpenStudy (anonymous):